Problem 7 A study of 427 college students was conducted to test whether high school GPA is a predictor of first-year college GPA. Students with higher High school GPA are expected to do better in college colgpa Grade point average in college (Range 0.85 3.97) hsgpa High school GPA (Range 2.29-4.5) Model 1: OLS, N -427 Dependent variable: colgpa coefficient 0.9 0.4 std. error const hsgpa 0.15 R-squared: 0.854880 a. (3%) Write the equation for the least-squares regression line: y- b. (3%) The nul hypothesis is: β Coonaa。 c (396) Calculate the t-ratio- d.。%)The critical value for the t distribution ( α-005 anddf-N-2 )s e. (3%) we reject the null hypothess? epok one
Answer #1

a)Yhat=0.9+0.4x

b)null hypothesis is :\beta <=0

c) t ratio =coefficient/std error =0.4/0.15=2.67

d)critical value =1.648

e) Yes (as test statistic is higher then critical value)